Explain the instruction format in detail as per the no of address in the instruction with examples?


1
Expert's answer
2021-11-03T04:39:02-0400

Instruction size = 32-bit


Number of instruction n = 98


a. number of bits for opcode

"ceil\left(\log{n}\right)"


"ceil\left(\log{98}\right)"


7 bits for opcode will be required


b.

number of bit in operand = (number of bit for instruction) - (bit required in opcode)

"32 - 7 = 25"

= 25 bits

c.

since there is only one Address and operand bit in instruction are 25


"\therefore maximum\ memory\ address\ that\ can\ be \ can\ be\ accessed \ is\ 2^{25}"


d.

yes it is because

In 7 bit opcode "2^7" = 128 instruction possible, but only 98 instructions are used in this set.


Total number of instruction may additionally use = "\left(128\ -\ 98\right)" "= 30"